# Language models scale reliably with over-training and on downstream tasks

### 13 Mar 2024 v1

@@ 3.1 Training setup
Architecture.
We train transformers [114], based on auto-regressive, decoder-only, pre-normalization architectures like GPT-2 [83] and LLaMA [111]. We adopt OpenLM [37] as our core modeling library, which utilizes PyTorch [78, 6], xformers [52], triton [73], FlashAttention [22], FSDP [127], and bfloat16 automatic mixed precision. Like LLaMA, we omit bias terms, but replace RMSNorm [125] with LayerNorm [8], which has readily available fused implementations. Following Wortsman et al. [120], we apply qk-LayerNorm [23], which adds robustness to otherwise poor hyperparameter choices (e.g., for learning rate). We use SwiGLU [100] activations and depth-scaled initialization [126]. We use a sequence length of 2048, rotary positional embeddings [104], and the GPT-NeoX-20B tokenizer [13], which yields a vocabulary size of 50k. We do not use weight tying [82, 44].
Objectives and optimization.
@@
Training datasets.
To ensure our conclusions are not particular to a training distribution, we train models on C4 [86, 25], RedPajama [110], and RefinedWeb [80]. They are open-source and have 138B, 1.15T, and 600B tokens respectively. We sample without replacement and employ sequence packing without attention masking. We separate documents in our training corpora with end-of-text tokens.
3.2 Creating scaling laws for validation loss and downstream error prediction
Figure 4: Search, filter, fit: A recipe for selecting configurations for scaling. (left) To generate the final configurations presented in Table 1, we run a 435 model grid search over model width, hidden dimension, number of attention heads, batch size, and warmup steps. All models are trained near compute-optimally. (center) We plot the efficient frontier of models, which appear to follow a trend, excluding models from $5.2\times 10^{16}$ to $5.2\times 10^{17}$, which fall below the trend. (right) We fit a power law with irreducible error to the remaining configurations, picking four configurations that closely track the full model suite. These models extrapolate the performance of 1.4B, 6.9B target models. Shaded regions represent bootstrap 95% confidence intervals.
Recipe for model configuration selection.
To create a testbed of models for our scaling experiments, we grid search over a total of 435 models, trained from scratch, in the 0.01B to 0.5B parameter range as seen in Figure 4 (left). We train on the OpenLM data mix [37], which largely consists of tokens from RedPajama [110] and The Pile [29]. We train on 20 tokens per parameter ($M=20$), which we find in early experiments gives models near the compute-optimal frontier for the data mix. This is similar to findings presented in Hoffmann et al. [43]’s Table 3, which suggests that roughly 20 tokens per parameter are optimal in their experimental setup. Our validation set, OpenLM eval, contains tokens from recent arXiv papers, the OpenLM codebase itself, and news articles. To find maximally performant models on validation data, we tune model width, number of layers, number of attention heads, warmup steps, and batch size. We find in early experiments that qk-LayerNorm makes models less sensitive to learning rate, which is a phenomenon Wortsman et al. [120] report in their Figure 1. Hence, we fix the learning rate for our sweeps. We also perform smaller grid searches over 1.4B and 6.9B parameter model configurations at $M=20$, retaining the best configurations.
@@ 3.2 Creating scaling laws for validation loss and downstream error prediction
In Figure 4 (center), we plot the efficient frontier of minimum loss configurations. While there appears to be a general trend, configurations between $5.2\times 10^{16}$ and $5.2\times 10^{17}$ FLOPs lie below the frontier established by other models. We hypothesize these models over-perform as they are trained for more optimization steps than their neighbors based on our power-of-two batch sizes. We provide support for this hypothesis in Appendix E, but opt to remove these models from our investigation.
In Figure 4 (right), we fit trends to the remaining models and to a subset of four models. We notice that the trends hit both the 1.4B and 6.9B models suggesting that our small-scale configurations are reasonable for extrapolation to larger parameter and compute regimes. We retain the four model configuration subsets as a representative sample.
@@
We do not tune hyperparameters for other token multipliers (i.e., $M\neq 20$), other training or evaluation distributions, or on downstream task validation sets. For more details, see Appendix C. We present our final hyperparameters in Table 1 given their importance.
[…]
Table 1: Main models and hyperparameters used in our investigation. Models have number of parameters $N$, with number of layers $n_{\text{layers}}$, number of attention heads $n_{\text{heads}}$, model width $d_{\text{model}}$, and width per attention head $d_{\text{head}}$. Batch sizes are global and in units of sequences. Each sequence has 2048 tokens. A100 GPU hours are at $M=20$, which are near compute-optimal runs. For the 1.4B scale, a batch size of 256 performs slightly better than 512.
Fitting scaling laws.
@@
Table 2: Default $N,M$ to fit our scaling laws. We invest $\sim$100 A100 hours to fit Equation (4) and $\sim$1000 A100 hours to fit Equation (5).
We fit Equation (4) to approximate $E,a,b,\alpha_{C}$ using curve-fitting in SciPy [115] (i.e., Levenberg-Marquardt to minimize non-linear least squares). We try several initializations and retain the best fit. We repeat this process to fit Equation (5) to approximate $\epsilon,k,\gamma$. Unless otherwise specified, we fit to the $N,M$ pairs in Table 2. In total, we invest $\sim$100 A100 hours to train the models required for fitting an accurate scaling law for loss prediction and $\sim$1000 A100 hours for a corresponding scaling law for downstream error prediction. Our configurations allow us to test for extrapolation to the $N=1.4\text{B},M=640$ (900B token) and the $N=6.9\text{B},M=20$ (138B token) regimes.
@@ 3.3 Evaluation setup
Unless otherwise stated, C4 eval is our default validation loss dataset. For downstream tasks, we adopt 46 tasks from LLM-foundry [67], which includes standard tasks with both zero-shot and few-shot evaluations. We also consider a 17-task subset where, for each evaluation, at least one 0.154B scale model—trained with as many as 99B tokens—gets 10 percentage points above chance accuracy: ARC-Easy [21], BIG-bench: CS algorithms [11], BIG-bench: Dyck languages [11], BIG-bench: Novel Concepts [11], BIG-bench: Operators [11], BIG-bench: QA WikiData [11], BoolQ [19], Commonsense QA [105], COPA [90], CoQA [89], HellaSwag (zero-shot) [123], HellaSwag (10-shot) [123], LAMBADA [75], PIQA [12], PubMed QA Labeled [48], SQuAD [88], and WinoGrand [53]. This subset allows us to investigate a regime where performance for small models can be non-trivial. For more details on evaluation datasets see Appendix D. For ablations on our choices of loss and downstream evaluations see Appendix E.
@@
We consider three main metrics: (i) Validation loss, which is the cross entropy between a model’s output and the one-hot ground truth, averaged over all tokens in a sequence and over all sequences in a dataset. (ii) Average top-1 error, which is a uniform average over 46 downstream evaluations sourced from LLM-foundry [67]. We also look at the mean top-1 error for the subset of 17 evaluations identified in the paragraph above. For a complete list of downstream evaluation datasets, see Appendix D. To measure how good a prediction $\zeta(C,M)$ is, we measure (iii) Relative prediction error: $|\zeta(C,M)-\zeta_{GT}|/\zeta_{GT}$, where $\zeta$ is the loss $L$ or the average top-1 error Err.
Testbed.
We train models on C4, RedPajama, and RefinedWeb with the number of parameters $N\in\{0.011\text{B},0.079\text{B},0.154\text{B},0.411\text{B}\}$ and token multipliers $M\in\{5,10,20,40,80,160,320,640\}$. We omit runs that require more tokens than are present in a dataset (i.e., $N=0.411\text{B},M=640$ for C4). We additionally train $N=1.4$B models at $M=20$ and at the largest token multiplier possible without repeating tokens (i.e., 80 for C4, 640 for RedPajama, and 320 for RefinedWeb). We train $N=6.9\text{B},M=20$ for each dataset. In total this results in 104 models. We evaluate each model on C4 eval for validation loss and on the 46 downstream tasks for top-1 error.
4 Results
Unless otherwise stated, we fit Equations (4), (5) to the configurations in Table 2 and use C4 eval for loss computation.
@@ 4 Results
These results support several key takeaways. (i) scaling can be predictable even when one increases the model size and the amount of over-training compared to the training runs used to fit a scaling law. (ii) Scaling can be predictable even in the presence of a distribution shift (e.g., RedPajama training and C4 evaluation). (iii) The form presented in Equation (4) is useful in practice for fitting and predicting over-trained scaling behavior. (iv) Fitting to Equation (4) does not sacrifice prediction accuracy near compute-optimal.
Figure 5: Relative error on C4 eval for different training distributions. Boxes highlighted in yellow correspond to data points used to fit Equation (4), all other points are predicted based on the fit. The prediction error is low in both interpolation and extrapolation ranges for number of parameters $N$ and token multiplier $M$, with larger values of $M$ specifying more over-training. Empty squares correspond to runs that were not possible due to compute limitations at the 1.4B and 6.9B scales or to limited dataset size for single epoch training.
While Figure 1 explores a specific case of making predictions in the over-trained regime, we would like to understand the error profile of our predictions across datasets, token multipliers, and number of parameters. Hence, in Figure 5 we show relative error between ground truth loss and predicted loss on C4 eval for models in our testbed. We notice uniformly low prediction error suggesting that predictions are accurate in many settings.
@@
Figure 1 (right) presents our results in estimating scaling laws for downstream error. Similar to Figure 1 (left), we are able to extrapolate in $N,M$ and predict the average downstream error across our evaluation suite. Concretely, we use the models indicated in Table 2 to fit Equation (5), and predict the average top-1 error over the 17 tasks identified in Section 3.3. Our fit allows us to predict the downstream performance of a 6.9B parameter model trained on 138B tokens to within $0.05\%$ relative error and of a 1.4B model trained on 900B tokens to within $3.6\%$ relative error, using 20×20$\times$20 × less compute.
@@
Table 3 additionally shows the relative error of our downstream performance predictions for models trained on C4, RedPajama, and RefinedWeb, indicating that our scaling law fits are applicable to other datasets. We note that while average accuracy across benchmarks is predictable, predicting accuracy on individual downstream tasks is significantly more noisy. We report relative error of all our predictions in Figures 11, 12 in the Appendix. We also find that if we remove the 1.4B model for the Equation (5) fit, relative error jumps, for instance, from 0.05% to 10.64% on the 17-task split for the 6.9B, 138B token RedPajama prediction.
Individual top-1 error Average top-1 error
Train set MMLU [38] OpenBook QA [66] HellaSwag [123] 17-task split 46-task split
C4 [86, 25] 2.82% 16.80% 79.58% 0.42% 0.76%
RedPajama [110] 0.12% 8.44% 25.73% 0.05% 2.10%
RefinedWeb [80] 0.77% 1.92% 81.96% 2.94% 2.76%
@@
Table 3: Downstream relative prediction error at 6.9B, 138B tokens. While predicting accuracy on individual zero-shot downstream evaluations can be challenging (“Individual”), predicting averages across downstream datasets is accurate (“Average”). On the right, we report relative prediction error for the average over a 17-task subset and over the full 46-task suite.
Small-scale experiments can predict model rank order.
We expect to be able to rank models based on their predicted performance, which is useful when deciding what to train. To verify, we rank 9 testbed models with $N\geq 1.4\text{B}$ by ground-truth top-1 error and by estimated top-1 error. We find high rank correlations: 0.93 and 0.88 for the 46 and 17-task splits respectively.
@@
Under-training, out-of-distribution scaling, and compute-reliability trade-offs.
@@
In addition to our main results presented above, we include additional results in Appendix E, which we summarize here. First, we notice that when token multipliers become too small (i.e., $M=5$) scaling becomes unreliable and lies off the trend. Additionally, several multipliers (10, 20, 40, and 80) garner points that are roughly on the compute optimal frontier (Figure 9). To probe the limits of reliable scaling, we attempt to break our scaling laws in out-of-distribution settings. We find that models trained on C4—English filtered—and evaluated on next token prediction on code domains have a high relative error in many cases. Perhaps surprisingly, evaluating the same models on German next token prediction again gives reliable loss scaling (Figure 10). We additionally examine the compute necessary to create accurate scaling laws, finding a positive correlation between investing more compute in a scaling law and its predictivity. We find that scaling laws can be constructed more cheaply for loss prediction than for downstream error prediction (Figures 1516).
@@ 5 Related work
We review the most closely related work in this section. For additional related work, see Appendix F.
Model family Parameters $N$ Training tokens $D$ Token multiplier $M$
T5 [87] 11B 34B 3.1
GPT-3 [14] 175B 300B 1.7
Gopher [84] 280B 300B 1.1
Chinchilla [43] 70B 1.4T 20.0
LLaMA [111] 7B 1T 140.0
LLaMA [111] 70B 1.4T 20.0
LLaMA-2 [112] 7B 2T 290.0
LLaMA-2 [112] 70B 2T 30.0
XGen [72] 7B 1.5T 210.0
MPT [108] 7B 1T 140.0
Table 4: Token multipliers of existing models. In our work, we run experiments with token multipliers between 5 and 640 for (GPT-2 [83], LLaMA [111])-style decoder-only architectures.
@@
Early works on scaling artificial neural networks observe predictable power-law scaling in the training set size and number of model parameters [41, 42, 91]. Alabdulmohsin et al. [2] center the importance of looking at the extrapolation regime of a scaling law. Yang et al. [121] prescribe architectural and hyperparameter changes when scaling model width to realize performant models; Yang et al. [122] make analogous recommendations when scaling model depth. Unlike the aforementioned work, our investigation focuses on the link between over-training models and predicting their downstream performance on accuracy metrics.
@@
Hoffmann et al. [43] investigate how the number of model parameters $N$ and training tokens $D$ should be chosen to minimize loss $L$ given a compute budget $C$. Hoffmann et al. [43] find that when scaling up $C$, both $N$ and $D$ should be scaled equally up to a multiplicative constant (i.e., $N\propto C^{\sim 0.5}$ and $D\propto C^{\sim 0.5}$) to realize compute-optimality. Appendix C of the Chinchilla paper additionally suggests that these findings hold across many datasets. However, Hoffmann et al. [43] do not account for inference costs, provide scaling laws for training beyond compute-optimal, or for downstream error prediction—all of which are central to our work.
@@
Sardana & Frankle [96] proposed modifications to the Chinchilla formulation to incorporate inference costs into the definition of compute-optimal and solve for various fixed inference budgets. Their key finding, which is critical for our work, is that when taking into account a large enough inference budget, it is optimal to train smaller models for longer when compared to the original Chinchilla recommendations. Our work presupposes that over-training can be beneficial. Instead of solving for inference-optimal schemes, we support empirically a predictive theory of scaling in the over-trained regime. Additionally, we provide experiments across many validation and training sets.
@@
For predicting downstream scaling beyond loss, Isik et al. [45] relate the number of pre-training tokens to downstream cross-entropy and machine translation BLEU score [76] after fine-tuning. In contrast, we do not examine a specific domain but rather take a holistic approach to evaluation by looking at top-1 error over many natural language tasks. Schaeffer et al. [98] argue that emergent abilities [118] and unreliable scaling are a product of non-linear metrics like error/accuracy and propose smoother alternatives. As an explanation for why non-linear metrics may be hard to predict, Schaeffer et al. [98] consider predicting exactly an $\ell$ length sequence: $\textsf{Err}(N,\ell)\approx 1-PP(N)^{-\ell}$, where $N$ is the number of parameters in a model and $PP$ is its perplexity. This is a special case of our Equations (5), (6), where number of training tokens is ignored, $\epsilon=1,k=1$, and $\gamma=\ell$. In contrast, we treat $\epsilon,k,\gamma$ as free parameters for a scaling law fit and embrace top-1 error, finding that average error over downstream tasks can make for a predictable metric.
@@
There has been a rise in over-trained models [111, 112] and accompanying massive datasets [110, 80, 102, 3]. To contextualize the extent to which we over-train, we provide token multipliers for popular models in Table 4. For example, Chinchilla 70B [43] is trained with a token multiplier of 20, while LLaMA-2 7B [112] uses a token multiplier of 290. In our investigation, we look at token multipliers from 5 to 640 to ensure coverage of popular models and relevance for future models that may be trained on even more tokens.

### 14 Jun 2024 v2

@@ 3.1 Training setup
We train transformers [116] for next token prediction, based on architectures like GPT-2 [85] and LLaMA [113]. We employ GPT-NeoX [15] as a standardized tokenizer for all data. See Appendix C for architecture, optimization, and hyperparameter details.
3.2 Model configurations
To get final configurations for the 0.011B to 0.411B parameter models plotted in Figures 2 and 3, we first conduct a wide grid search over a total of 435 models, trained from scratch, from 0.01B to 0.5B parameters (Figure 4 (left)). We train on the original OpenLM data mix [39], which largely consists of RedPajama [112] and The Pile [31]. While we eventually plan to over-train models, at this step we search for base configurations near compute-optimality. We train on 20 tokens per parameter ($M=20$), which, in early experiments, gives models near the compute-optimal frontier. This is similar to findings in Hoffmann et al. [45]’s Table 3, which suggests that $M=20$ is near-optimal for the Chinchilla experimental setup.
@@ 3.2 Model configurations
At this point, we have many models, several of which give poor performance; following prior work [51, 45], we want to keep only models that give best performance. Hence, in Figure 4 (center), we filter out models that do not lie on the Pareto frontier. While there appears to be a general trend, configurations between $5.2\times 10^{16}$ and $5.2\times 10^{17}$ FLOPs lie below the frontier established by other models. We hypothesize these models over-perform as they are trained for more optimization steps than their neighbors based on our power-of-two batch sizes. We provide support for this hypothesis in Appendix F, but opt to remove these models from our investigation.
To ensure tractable compute requirements for our scaling experiments, we require a subset of models that follows the trend of the entire Pareto frontier. In Figure 4 (right), we fit trends to the Pareto models and to a subset of four models. We notice that the trends closely predict both the performance of the 1.4B and 6.9B models, suggesting that our small-scale configurations reliably extrapolate in the compute-optimal setting.
@@
Moving forward, we do not tune hyperparameters for other token multipliers (i.e., $M\neq 20$), on other training or evaluation distributions, or on validation sets for downstream tasks. For more details including specific hyperparameters, see Appendix D.
To create our scaling testbed, we start with the four small-scale, base configurations from our grid search: $N\in\{0.011\text{B},0.079\text{B},0.154\text{B},0.411\text{B}\}$. To ensure our conclusions are not particular to a single training distribution, we train models on each of C4 [88, 27], RedPajama [112], and RefinedWeb [82], which have 138B, 1.15T, and 600B tokens, respectively, for different token multipliers $M\in\{5,10,20,40,80,160,320,640\}$. We omit runs that require more tokens than are present in a dataset (i.e., $N=0.411\text{B},M=640$ for C4). We additionally train $N=1.4$B models at $M=20$ and at the largest token multiplier possible without repeating tokens (i.e., 80 for C4, 640 for RedPajama, and 320 for RefinedWeb). We train $N=6.9\text{B},M=20$ models on each dataset given the relevance of 7B parameter models [113, 49]. In total this results in a testbed of 104 models.
3.3 Fitting scaling laws
Table 1: Default number of parameters $N$ and token multiplier $M$ to fit our scaling laws. We invest $\sim$100 A100 hours to fit Equation (4) and $\sim$1,000 A100 hours to fit Equation (5).
@@ 3.3 Fitting scaling laws
We fit Equation (4) to approximate $E,a,b,\eta$ using curve-fitting in SciPy [117] (i.e., Levenberg-Marquardt to minimize non-linear least squares). We repeat this process to fit Equation (5) to approximate $\epsilon,k,\gamma$. We invest $\sim$100 A100 hours to train the models required to fit a scaling law for loss and $\sim$1,000 A100 hours for a corresponding law for downstream error. Unless otherwise specified, we fit to the $N,M$ pairs in Table 1, which are a subset of our full testbed. Our configurations allow us to test for extrapolation to the $N=1.4\text{B},M=640$ (900B token) and the $N=6.9\text{B},M=20$ (138B token) regimes.
@@ 3.4 Evaluation setup
Unless otherwise stated, our default validation loss dataset is C4 eval. For downstream tasks, we adopt a subset from 46 tasks from LLM-foundry [69], which includes standard tasks with both zero-shot and few-shot evaluations. Specifically, we consider a 17-task subset where, for each evaluation, at least one 0.154B scale model—trained with as many as 99B tokens—gets 10 percentage points above chance accuracy: ARC-Easy [23], BIG-bench: CS algorithms [11], BIG-bench: Dyck languages [11], BIG-bench: Novel Concepts [11], BIG-bench: Operators [11], BIG-bench: QA WikiData [11], BoolQ [21], Commonsense QA [107], COPA [92], CoQA [91], HellaSwag (zero-shot) [126], HellaSwag (10-shot) [126], LAMBADA [77], PIQA [14], PubMed QA Labeled [50], SQuAD [90], and WinoGrand [55]. For more details on evaluation datasets see Appendix E. We focus on this subset to ensure we are measuring signal, not noise. Including downstream tasks like MMLU [40], where performance is close to random chance, however, does not invalidate our results as we show in our evaluation set ablations (Appendix F).
@@
We consider three main metrics: Validation loss, which is the cross entropy between a model’s output and the one-hot ground truth token, averaged over all tokens in a sequence and over all sequences in a dataset. Average top-1 error, which is a uniform average over the 17 downstream evaluations, as mentioned in the above paragraph. To measure how good a prediction $\zeta(C,M)$ is, we measure Relative prediction error: $|\zeta(C,M)-\zeta_{GT}|/\zeta_{GT}$, where $\zeta$ is the predicted loss $L$ or the average top-1 error Err. $\zeta_{GT}$ is the ground truth measurement to predict.
4 Results: Reliable extrapolation
In this Section, we quantify the extent to which the scaling laws developed in Section 2 extrapolate larger model performance using the scaling testbed from Section 3. By default, we fit Equations (4) and (5) to the configurations in Table 1, use C4 eval for loss, and the 17-task split from Section 3.4 for average top-1 error.
Figure 5: Relative error on C4 eval for different training distributions. Boxes highlighted in yellow correspond to pairs—number of parameters $N$, token multiplier $M$—used to fit Equation (4). Larger values of $M$ correspond to more over-training. The prediction error is low in both interpolation and extrapolation ranges. Below $N=1.4$B, empty squares correspond to runs that were not possible due to the limited dataset size for single epoch training. At $N=1.4$B we run at $M=20$ and at the largest possible multiplier. At $N=6.9$B, we run at $M=20$.
@@ 4 Results: Reliable extrapolation
These results support several key takeaways. (i) Scaling can be predictable even when one increases both the model size and the amount of over-training compared to the training runs used to fit a scaling law. (ii) The form presented in Equation (4) is useful in practice for predicting over-trained scaling behavior. (iii) Fitting to Equation (4) gives good prediction accuracy near compute-optimal. More specifically, predictions are accurate both for the 1.4B over-trained model and the 6.7B compute-optimal model using a single scaling fit.
While Figure 1 explores a specific case of making predictions in the over-trained regime, we aim to understand the error profile of our predictions across training datasets, token multipliers, and number of parameters. Hence, Figure 5 shows the relative error between ground truth loss and predicted loss on C4 eval for models in our testbed. We notice uniformly low prediction error suggesting that predictions are accurate in many settings.
@@
Figure 1 (right) presents our main result in estimating scaling laws for downstream error. Concretely, we use the models indicated in Table 1 to fit Equations (4) and (5), chaining the scaling fits to predict the average top-1 error as a function of training compute $C$ and the token multiplier $M$. Our fits allow us to predict, using 20×20$\times$20 × less compute, the downstream performance of a 6.9B model trained on 138B RedPajama tokens to within $0.05\%$ relative error and a 1.4B model trained on RedPajama 900B tokens to within $3.6\%$ relative error.
@@
Table 2 additionally shows the relative error of our downstream performance predictions for models trained on C4, RedPajama, and RefinedWeb, indicating that our scaling law functional forms are applicable on many training datasets. We note that while average accuracy is predictable, individual downstream task predictions are significantly more noisy. We report relative error for more model predictions in Figures 11 and 12. We also find that if we remove the 1.4B model for the Equation (5) fit, relative error jumps, for instance, from 0.05% to 10.64% on the 17-task split for the 6.9B, 138B token RedPajama prediction. This highlights the importance of investing more compute when constructing scaling laws for downstream task prediction compared to loss prediction.
@@
Table 2: Downstream relative prediction error at 6.9B parameters and 138B tokens. While predicting accuracy on individual zero-shot downstream evaluations can be challenging (“Individual”), predicting averages across downstream datasets is accurate (“Avg.”).
Individual top-1 error Avg. top-1 error
Train set ARC-E [23] LAMBADA [77] OpenBook QA [68] HellaSwag [126] 17-task split
C4 [88, 27] 28.96% 15.01% 16.80% 79.58% 0.14%
RedPajama [112] 5.21% 14.39% 8.44% 25.73% 0.05%
RefinedWeb [82] 26.06% 16.55% 1.92% 81.96% 2.94%
@@
Under-training, out-of-distribution scaling, compute-reliability trade-offs.
@@
In addition to our main results presented above, we include additional results in Appendix F, which we summarize here. First, we notice that when token multipliers become too small (i.e., $M=5$) scaling becomes unreliable and lies off the trend. Additionally, multipliers other than 20, such as 10, 40, and 80, garner points that are roughly on the compute optimal frontier (Figure 9). This observation suggests that the compute-optimal multiplier may lie in a range rather than take a single value. To probe the limits of reliable scaling, we attempt to break our scaling laws in out-of-distribution settings. We find that models trained on C4—English filtered—and evaluated on next token prediction on code domains have a high relative error in many cases. Perhaps surprisingly, evaluating the same models on German next token prediction gives reliable loss scaling (Figure 10). We additionally examine the compute necessary to create accurate scaling laws, finding that scaling laws can be constructed more cheaply for loss prediction than for downstream error prediction (Figures 15 and 16).
@@ 5 Related work
We review the most closely related work in this section. For additional related work, see Appendix G.
@@
Early works on scaling artificial neural networks observe predictable power-law scaling in the training set size and number of model parameters [43, 44, 93]. Alabdulmohsin et al. [2] stress the importance of looking at the extrapolation regime of a scaling law. Yang et al. [124] prescribe architectural and hyperparameter changes when scaling model width to realize performant models; Yang et al. [125] make analogous recommendations when scaling model depth. Bi et al. [13] propose hyperparameter aware scaling laws. Unlike the aforementioned work, our investigation focuses on over-training and predicting downstream accuracy.
@@
Hoffmann et al. [45] investigate how the number of model parameters $N$ and training tokens $D$ should be chosen to minimize loss $L$ given a compute budget $C$. Hoffmann et al. [45] find that when scaling up $C$, both $N$ and $D$ should be scaled equally up to a multiplicative constant (i.e., $N\propto C^{\sim 0.5}$ and $D\propto C^{\sim 0.5}$) to realize compute-optimality. Appendix C of the Chinchilla paper additionally suggests that these findings hold across three datasets. However, Hoffmann et al. [45] do not verify their scaling laws for training beyond compute-optimality, or for downstream error prediction—both of which are central to our work.
@@
Sardana & Frankle [98] propose modifications to the Chinchilla formulation to incorporate inference costs into the definition of compute-optimality and solve for various fixed inference budgets. Their key finding, which is critical for our work, is that when taking into account a large enough inference budget, it is optimal to train smaller models for longer than the original Chinchilla recommendations. Our work presupposes that over-training can be beneficial. Instead of solving for inference-optimal schemes, we support empirically a predictive theory of scaling in the over-trained regime. Additionally, we provide experiments across many validation and training sets.
@@
For predicting downstream scaling beyond loss, Isik et al. [47] relate the number of pre-training tokens to downstream cross-entropy and machine translation BLEU score [78] after fine-tuning. In contrast, we take a holistic approach to evaluation by looking at top-1 error over many natural language tasks. Schaeffer et al. [100] argue that emergent abilities [120] are a product of non-linear metrics and propose smoother alternatives. As a warmup for why non-linear metrics may be hard to predict, Schaeffer et al. [100] consider predicting an $\ell$ length sequence exactly: $\textsf{Err}(N,\ell)\approx 1-\textsf{PP}(N)^{-\ell}$, where $N$ is the number of parameters in a model and PP is its perplexity. This is a special case of our Equations (5) and (6), where the number of training tokens does not appear, $\epsilon=1,k=1$, and $\gamma=\ell$. In contrast, we treat $\epsilon,k,\gamma$ as free parameters for a scaling law fit, finding that average error over downstream tasks can make for a predictable metric.
@@
There has been a rise in over-trained models [113, 114] and accompanying massive datasets [112, 82, 104, 3]. For example, Chinchilla 70B [45] is trained with a token multiplier of 20, while LLaMA-2 7B [114] uses a token multiplier of 290. In our investigation, we look at token multipliers from 5 to 640 to ensure coverage of popular models and relevance for future models that may be trained on even more tokens.